# Quantitative 2D propagation of smallness and control for 1D heat equations with power growth potentials

### v1

@@ Abstract.
We study the relation between propagation of smallness in the plane and control for heat equations. The former has been proved by Zhu [Zhu23] who showed how the value of solutions in some small set propagates to a larger domain. By reviewing his proof, we establish a quantitative version with the explicit dependence of parameters. Using this explicit version, we establish new exact null-controllability results of 1D heat equations with any nonnegative power growth potentials $V\in\mathcal{C}(\mathbb{R})$. As a key ingredient, new spectral inequalities are established. The control set $\Omega$ that we consider satisfy
@@ 1.1. Main results
where potential $V$ is a real-valued nonnegative continuous funtion, $h(t,x)\in L^{2}\left((0,T)\times\mathbb{R}\right)$, and $\Omega\subset\mathbb{R}$ is a given measurable set. Equation (1.1) is said to be exactly null-controllable (or simply null-controllable) from the set $\Omega$ in time $T>0$ if, for any intial datum $u_{0}\in L^{2}(\mathbb{R})$, there exists $h(t,x)\in L^{2}\left((0,T)\times\Omega\right)$ such that the mild solution of (1.1) satisfies $u(T)=0$.
@@
$V(x)\in\mathcal{C}(\mathbb{R})$ is a continuous nonnegative real-valued potential and there exist constants $c_{1}>0$, $c_{2}>0$, $c_{3}>0$ and $\beta_{2}\geq\beta_{1}>0$ such that
@@
where $(a)_{+}:=\min\left\{0,a\right\}$ and $\langle x\rangle:=(1+|x|^{2})^{\frac{1}{2}}$ denotes the Japanese bracket.
@@
    If we fix $s=0$ and $\tau=0$, the $(L,\gamma)$-thick set $\Omega$ of type $(\rho_{0},0)$ is just the usual definition of a thick set, that is, there exist new constants $L>0$ and $\gamma>0$ such that for all $x\in\mathbb{R}$, we have
    (1.5) $|\Omega\cap[x,x+L]|\geq\gamma L.$
@@
(1.6) $\rho(x):=\min\left\{\left(R\log\log\langle x\rangle-R^{-1}\right)^{\frac{1}{2}}\frac{1}{\langle x\rangle^{\frac{\beta_{2}}{2}}},1\right\}$
@@
$V\in\mathcal{C}(\mathbb{R})$ is a continuous nonnegative real-valued potential satisfing Assumption A1. We say that $V$ satisfies Assumption A2 if it can be written as $V=V_{1}+V_{2}$ and
@@
(1.9) $\|e^{-Ht}u_{0}\|^{2}_{L^{2}(\mathbb{R})}\leq C(T,V,\Omega)\int_{0}^{T}\|e^{-Ht}u_{0}\|^{2}_{L^{2}(\Omega)}\,\mathrm{d}t,\quad\forall u_{0}\in L^{2}(\mathbb{R}),$
@@ 1.2. Spectral inequalities
A spectral inequality for a nonnegative selfadjoint operator $P$ in $L^{2}(\mathbb{R}^{d})$ takes the form
@@
(1.16) $f(\sqrt{H})=\int_{0}^{\infty}F(\lambda)\,\mathrm{d}m_{\lambda},\quad\forall f\in L^{\infty}(\mathbb{R}),$
@@
(1.17) $(F(\sqrt{H})f,G(\sqrt{H})f)=\int_{0}^{\infty}F(\lambda)\overline{G(\lambda)}\left(\,\mathrm{d}m_{\lambda}f,f\right)_{L^{2}(\mathbb{R})},\quad\forall f\in L^{2}(\mathbb{R})$
for any $F,G\in L^{\infty}(\mathbb{R})$. Then the spectral projector $\Pi_{\mu}$ associated with the function $F(\lambda)=\mathbbm{1}_{\lambda\leq\mu}$, is defined by
(1.18) $\Pi_{\mu}f:=\int_{0}^{\mu}F(\lambda)\,\mathrm{d}m_{\lambda}.$
@@
    The exponential in Theorem 1.9 is $\lambda^{(4\tau+2)/\beta_{1}+\beta_{2}/\beta_{1}}+\lambda^{{\beta_{2}}/{\beta_{1}}}\log(\lambda+1)$, while the one in [ZZ23, Theorem 1] is $\lambda^{{2\tau}/{\beta_{1}}+{\beta_{2}}/{\beta_{1}}}$.
In [SSY23], the authors considered the observability inequality of the 1D Schrödinger equation for $V(x)\in\mathcal{C}(\mathbb{R})$ and bounded, and proved it for $\Omega$ being $\gamma$-thick. To do so, they established the spectral inequality for the Schrödinger operator $H=-\partial_{x}^{2}+V(x)$ with $V(x)\in\mathcal{C}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$, we present it here as a comparison:
@@
Let $\Omega$ be a $(1,\gamma)$-thick set and let $V\in\mathcal{C}(\mathbb{R})\cap L^{\infty}$. Then there exists a constant $C=C(V,\gamma)>0$ depending only on $V$ and $\gamma$, such that for any $\lambda>0$ and any $f\in L^{2}(\mathbb{R})$, we have
@@
$\|e^{-tP}g\|^{2}_{L^{2}(\mathbb{R}^{d})}\leq\frac{C_{\mathrm{obs}}}{T}\int_{0}^{T}\|e^{-tP}g\|^{2}_{L^{2}(\omega)}\,\mathrm{d}t,$
@@ 1.3. Previous results
There have been numerous studies on the null-controllability of heat equations over the past decades, both on bounded domains with various boundary conditions and on unbounded domains. As our results deal with the heat equation on $\mathbb{R}$, we will focus on the Euclidean case, or some results that are highly related to the propagation of smallness. For the results on bounded domain, we may refer e.g., to [LR95, LZ98, AE08, LRL12, AE13, Lü13, AEWZ14, EMZ17, BM22].
@@
For the harmonic case $V=|x|^{2}$, the spectral inequality (and hence the null-controllability of the heat equation) from different kinds of $\Omega$ was proved in [BPS18, BJPS21, ES21, MPS22, DSV23]. All of them are based on the explicit form of the Hermite functions which are the eigenfunctions of the harmonic oscillator. In particular, we mention that [DSV23] allows control sets $\Omega\subset\mathbb{R}^{d}$ satisfying
@@
For a less regular potential, all strategies mentioned above seem not to work. However, the spectral inequality for the 1D Schrödinger operator with bounded continuous potentials from thick sets was proved in [SSY23]. This relies on the new result about the propagation of smallness for elliptic equations in the plane, see [Zhu23]. The proof is limited to 1D because, in higher dimensions, we only have a weaker quantitative propagation of smallness, see [LM18] for the propagation of smallness in higher dimensions. Inspired by their work, we generalize their results to power growth potentials. The new obstacles occur in unbounded potentials, which will be discussed in the next subsection.
@@ 1.4. Outline of the proof
(1.29) $\nabla\cdot(A\nabla\Psi)=0,\quad A\nabla\Psi\cdot\mathbf{e}_{0}=0\text{ in }\ell_{0},$
@@ 2.1. Reduction of thick sets
Let $\left\{x_{n}\right\}_{n\in\mathbb{N}}$ be the sequence given by (2.1) for some $L>0$, then we have
@@
First of all, we obtain immediately from (2.1) that the sequence $\left\{x_{n}\right\}_{n\in\mathbb{N}}$ is increasing and $\lim_{n\to\infty}x_{n}=\infty$. What we need to prove is the relation $x_{n}^{s+1}\sim(s+1)Ln$. Indeed we have
@@
We define for all $n\in\mathbb{N}$
(2.8) $c_{n}:=\frac{x_{n}+x_{n+1}}{2}=x_{n}+\frac{L}{2}x_{n}^{-s}$
@@
(2.12) $|\Omega\cap I_{n}|\geq|\Omega\cap I_{L^{\prime}\rho_{s}(c_{n})}(c_{n})|\geq{\gamma^{\prime}}^{\langle c_{n}\rangle^{\tau}}|\Omega\cap I_{\rho_{s}(c_{n})}(c_{n})|=\frac{2{\gamma^{\prime}}^{\langle c_{n}\rangle^{\tau}}}{\langle c_{n}\rangle^{s}}.$
@@
(2.14) $|I_{n}|=x_{n}^{-s}\leq C_{1}^{-s}\langle c_{n}\rangle^{-s}.$
@@
(2.15) $|\Omega\cap I_{n}|\geq 2C_{1}^{s}{\gamma^{\prime}}^{\langle c_{n}\rangle^{\tau}}|I_{n}|.$
The inequality (2.6) is obtained from the above by choosing $\gamma$ small enough so that $\gamma<\inf_{n\in\mathbb{N}}(2C_{1}^{s})^{1/\langle c_{n}\rangle^{\tau}}\gamma^{\prime}$, where the right hand side is positive since $\langle c_{n}\rangle\geq 1$ for all $n\in\mathbb{N}$.
@@
(2.20) $L^{\prime}>\max_{k\in\mathbb{N}_{+}}\left\{2L\langle L\rangle^{s},\left(\frac{\langle x_{k+1}\rangle}{x_{k}}\right)^{s}\right\}\geq\max\left\{2L\langle L\rangle^{s},\left(\frac{\langle x_{1}\rangle}{x_{0}}\right)^{s}\right\},$

### v2

@@ Abstract.
We study the relation between propagation of smallness in the plane and control for heat equations. The former has been proved by Zhu [Zhu23] who showed how the value of solutions in some small set propagates to a larger domain. By reviewing his proof, we establish a quantitative version with the explicit dependence of parameters. Using this explicit version, we establish new exact null-controllability results of 1D heat equations with any nonnegative power growth potentials $V\in L^{\infty}_{\mathrm{loc}}(\mathbb{R})$. As a key ingredient, new spectral inequalities are established. The control set $\Omega$ that we consider satisfy
@@ 1.1. Main results
where potential $V$ is a real-valued nonnegative funtion, $h(t,x)\in L^{2}\left((0,T)\times\mathbb{R}\right)$, and $\Omega\subset\mathbb{R}$ is a given measurable set. Equation (1.1) is said to be exactly null-controllable (or simply null-controllable) from the set $\Omega$ in time $T>0$ if, for any intial datum $u_{0}\in L^{2}(\mathbb{R})$, there exists $h(t,x)\in L^{2}\left((0,T)\times\Omega\right)$ such that the mild solution of (1.1) satisfies $u(T)=0$.
@@
$V(x)\in L^{\infty}_{\mathrm{loc}}(\mathbb{R})$ is a nonnegative real-valued potential and there exist constants $c_{1}>0$, $c_{2}>0$, $c_{3}>0$ and $\beta_{2}\geq\beta_{1}>0$ such that
@@
where $(a)_{+}:=\max\left\{0,a\right\}$ and $\langle x\rangle:=(1+|x|^{2})^{\frac{1}{2}}$ denotes the Japanese bracket.
@@
    If we fix $s=0$, the $(L,\gamma)$-thick set $\Omega$ of type $(\rho_{0},\tau)$ is reduced to such a condition: there exist new constants $L>0$ and $\gamma>0$ and the same $\tau$ such that
    (1.5) $|\Omega\cap[x,x+L]|\geq\gamma^{\tau}L,\quad\forall x\in\mathbb{R}.$
    If we fix $\tau=0$ further, the $(L,\gamma)$-thick set $\Omega$ of type $(\rho_{0},0)$ is just the usual definition of a thick set, that is, there exist new constants $L>0$ and $\gamma>0$ such that
    (1.6) $|\Omega\cap[x,x+L]|\geq\gamma L,\quad\forall x\in\mathbb{R}.$
@@
(1.7) $\rho(x):=\min\left\{\left(R\log\log\langle x\rangle-R^{-1}\right)\frac{1}{\langle x\rangle^{\frac{\beta_{2}}{2}}},1\right\}$
@@
$V\in L^{\infty}_{\mathrm{loc}}(\mathbb{R})$ is a nonnegative real-valued potential satisfing Assumption A1. We say that $V$ satisfies Assumption A2 if it can be written as $V=V_{1}+V_{2}$ and
@@
(1.10) $\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}e^{-HT}}u_{0}\|^{2}_{L^{2}(\mathbb{R})}\leq C(T,V,\Omega)\int_{0}^{T}\|e^{-Ht}u_{0}\|^{2}_{L^{2}(\Omega)}\,\mathrm{d}t,\quad\forall u_{0}\in L^{2}(\mathbb{R}),$
@@ 1.2. Spectral inequalities
Given any $d$-dimensional Euclidean space $\mathbb{R}^{d}$, a spectral inequality for a nonnegative selfadjoint operator $P$ in $L^{2}(\mathbb{R}^{d})$ takes the form
@@
(1.17) $f=\int_{0}^{\infty}\,\mathrm{d}m_{\lambda}f,\quad\forall f\in L^{2}(\mathbb{R}).$
Moreover, we have
(1.18) $F(\sqrt{H})=\int_{0}^{\infty}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}F}(\lambda)\,\mathrm{d}m_{\lambda},\quad\forall F\in L^{\infty}(\mathbb{R}),$
@@
(1.19) $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G}(\sqrt{H})f,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}H}(\sqrt{H})f)=\int_{0}^{\infty}G(\lambda)\overline{H(\lambda)}\left(\,\mathrm{d}m_{\lambda}f,f\right)_{L^{2}(\mathbb{R})},\quad\forall f\in L^{2}(\mathbb{R})$
for any $G,H\in L^{\infty}(\mathbb{R})$. Then the spectral projector $\Pi_{\mu}$ associated with the function $F(\lambda)=\mathbbm{1}_{\lambda\leq\mu}$, is defined by
(1.20) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Pi_{\mu}f:=\int_{0}^{\mu}\,\mathrm{d}m_{\lambda}f,\quad\forall f\in L^{2}(\mathbb{R}).}$
@@
    The exponent in Theorem 1.9 is ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda^{4\tau/\beta_{1}+\beta_{2}/\beta_{1}}\log(\lambda+1)}$, while the one in [ZZ23, Theorem 1] is $\lambda^{{2\tau}/{\beta_{1}}+{\beta_{2}}/{\beta_{1}}}$. For a possible improvement of the exponent, there are two parts which are worse than the one in [ZZ23]. First, the extra $\log$ term arises from the application of Proposition 4.9, which is a spectral inequality for a more regular set $\widetilde{\Omega}$. This set includes a union of open intervals: there exists a sequence $\{x_{k}\}_{k\in\mathbb{Z}}\subset\mathbb{R}$ such that
    (1.24) $\exists L>0,\sigma\geq 0:\quad\widetilde{\Omega}\cap(k+[0,L])\supset I_{\langle k\rangle^{-\sigma}L}(x_{i}),\quad\forall k\in\mathbb{Z}.$
    We use Proposition 4.9 for some $\sigma>0$, and this will give the $\log$ term. Hence the $\log$ term cannot be dropped unless we do not use this auxilary result. It is an ad-hoc method to reduce our problem to the regular one by the pigeonhole principle so that we can use Carleman estimates. Second, even without the $\log$ term, it is still worse than the one in [ZZ23]. It relys on the estimates of the constants in Theorem 3.1. It is possible to give a sharper exponent if: (1) one can improve the estimate (3.4) in Theorem 3.1 by replacing $\bigl{|}\log|\omega|\bigr{|}^{2}$ with $\bigl{|}\log|\omega|\bigr{|}$; (2) one assumes extra regularity of potentials and use a totally different propagation of smallness property, see e.g., [Zhu24, Theorem 1].
Remark 1.11.
Compared to Theorem 1.7 and Theorem 1.8, in Theorem 1.9 we extend the set $\Omega$ from thick of type $(\rho_{s},\tau)$ to an $\Omega$ that is thick of type $(\rho_{0},\tau)$, giving $I_{L\rho(x)}=[x-L.x+L]$. We believe that it is possible to extend this result further to the case where $\rho(x)$ grows with $|x|\to\infty$. This has been done in [AS22, MPS23] for certain analytic potentials. For a general potential, it is not known whether we can extend it further under Assumption A2. Noticing that an extra regularity condition in Assumption A2 (compared to Assumption A1) is needed when we extend to $\rho_{0}$ case, therefore such an extension may needs some improvements of the regularity posed on the potential so that we have some sort of smoothing effects like those in [AS22, MPS23].
In [SSY23], the authors considered the observability inequality of the 1D Schrödinger equation for $V(x)\in L^{\infty}(\mathbb{R})$, and proved it for $\Omega$ being $\gamma$-thick. To do so, they established the spectral inequality for the Schrödinger operator $H=-\partial_{x}^{2}+V(x)$ with $V(x)\in L^{\infty}(\mathbb{R})$, we present it here as a comparison:
@@
Let $\Omega$ be a $(1,\gamma)$-thick set and let $V\in L^{\infty}(\mathbb{R})$. Then there exists a constant $C=C(V,\gamma)>0$ depending only on $V$ and $\gamma$, such that for any $\lambda>0$ and any $f\in L^{2}(\mathbb{R})$, we have
@@
$\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}e^{-PT}}g\|^{2}_{L^{2}(\mathbb{R}^{d})}\leq\frac{C_{\mathrm{obs}}}{T}\int_{0}^{T}\|e^{-tP}g\|^{2}_{L^{2}(\omega)}\,\mathrm{d}t,$
@@ 1.3. Previous results
There have been numerous studies on the null-controllability of heat equations over the past decades, both on bounded domains with various boundary conditions and on unbounded domains. As our results deal with the heat equation on $\mathbb{R}$, we will focus on the Euclidean case, or some results that are highly related to the propagation of smallness. We may refer e.g., to [LR95, LZ98, AE08, LRL12, AE13, Lü13, AEWZ14, EMZ17, BM22] for the results on bounded domains, and [Rou24] for the spectral inequality on the hyperbolic half-plane. Throughout the article, the use of $\mathbb{R}^{d}$ means that it works for any $d$-dimensional Euclidean space.
@@
For the harmonic case $V=|x|^{2}$, the spectral inequality (and hence the null-controllability of the heat equation) from different kinds of $\Omega$ was proved in [BPS18, BJPS21, ES21, MPS23, DSV23]. All of them are based on the explicit form of the Hermite functions which are the eigenfunctions of the harmonic oscillator. In particular, we mention that [DSV23] allows control sets $\Omega\subset\mathbb{R}^{d}$ satisfying
@@
For a less regular potential, all strategies mentioned above seem not to work. However, the spectral inequality for the 1D Schrödinger operator with bounded potentials from thick sets was proved in [SSY23]. This relies on the new result about the propagation of smallness for elliptic equations in the plane, see [Zhu23]. The proof is limited to 1D because, in higher dimensions, we only have a weaker quantitative propagation of smallness, see [LM18] for the propagation of smallness in higher dimensions. Inspired by their work, we generalize their results to power growth potentials. The new obstacles occur in unbounded potentials, which will be discussed in the next subsection.
Unlike the case $V=0$, we have very limited knowledge about some kinds of necessary conditions on the observable sets for the Schrödinger operators power growth potentials. One possible way is to use the same strategy in the proof of the case $V=0$ in [WWZZ19]. This needs the pointwise estimate of the heat kernel for the Schrödinger operator with the power growth potential, see e.g., an estimate in [Sik97]. Besides, we need to choose a suitable way to adapt the proof in [WWZZ19]. We do not pursue this direction here.
1.4. Latest results on arbitrary dimensions
In the context of arbitrary $\mathbb{R}^{d}$ or compact manifolds, there are two very recent preprints [Zhu24, LM24] that were posted on the arxiv shortly after this paper. They contain results for both unbounded and bounded rough potentials that are closely related to this work:
In [Zhu24], the author considered power growth potential $V\in L_{\mathrm{loc}}^{\infty}(\mathbb{R}^{d})\cap\mathcal{C}^{0,1}_{\mathrm{loc}}(\mathbb{R}^{d}\backslash\mathcal{B}_{R})$ for some $R>0$, and established Theorem 1.9 with an exponent as sharp as the one in [ZZ23]. He used a different way to construct an extra ghost dimension to absorb the potential into the principal part of the operator. The author also considered the bounded potential and established Theorem 1.12 in this case as well as a new quantitative propagation of smallness for gradients of solutions to some elliptic equations (see [Zhu24, Lemma 4]). Nevertheless, this work requires extra regularity condition $V\in\mathcal{C}^{0,1}_{\mathrm{loc}}(\mathbb{R}^{d}\backslash\mathcal{B}_{R})$ on power growth potentials to obtain a sharper exponent.
In [LM24], the authors considered $V\in L^{\infty}(M)$ both when $M$ is a compact manifold and when $M=\mathbb{R}^{d}$. They use an auxiliary function (see [LM24, Lemma 3.1]) to absorb the potential and so reducing the operator to a divergence form one. Based on two propagation of smallness results (see [LM18, Theorem 2.1, Theorem 5.1]), they obtain new propagation of smallness properties for nondivergence form operators (see [LM24, Theorem 2.1, Theorem 2.2]).
@@ 1.5. Outline of the proof
(1.32) $\nabla\cdot(A\nabla\Psi)=0{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{ in }B_{4}},\quad A\nabla\Psi\cdot\mathbf{e}_{0}=0\text{ in }\ell_{0},$
@@ 2.1. Reduction of thick sets
Let ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left\{x_{n}\right\}_{n\in\mathbb{N}_{+}}}$ be the sequence given by (2.1) for some $L>0$, then we have
@@
First of all, we obtain immediately from (2.1) that the sequence ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left\{x_{n}\right\}_{n\in\mathbb{N}_{+}}}$ is increasing and $\lim_{n\to\infty}x_{n}=\infty$. Indeed, if $x_{n}\leq C$ for some positive constant $C>0$, then
(2.5) $x_{n+m}-x_{n}=\sum_{k=0}^{m-1}\bigl{(}x_{n+k}-x_{n+k}\bigr{)}\geq m\frac{1}{C^{s}}\to\infty\text{ when }m\to\infty,$
which makes a contradiction. What we need to prove is the relation $x_{n}^{s+1}\sim(s+1)Ln$. Indeed we have
@@
We define
(2.9) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}c_{0}=\frac{L}{2}\text{ and }c_{n}:=\frac{x_{n}+x_{n+1}}{2}=x_{n}+\frac{L}{2}x_{n}^{-s},\quad\forall n\in\mathbb{N}_{+}.}$
@@
(2.13) $|\Omega\cap I_{n}|\geq|\Omega\cap I_{L^{\prime}\rho_{s}(c_{n})}(c_{n})|\geq{\gamma^{\prime}}^{\langle c_{n}\rangle^{\tau}}|\Omega\cap I_{\rho_{s}(c_{n})}(c_{n})|={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{2L^{\prime}{\gamma^{\prime}}^{\langle c_{n}\rangle^{\tau}}}{\langle c_{n}\rangle^{s}}}.$
@@
(2.15) $|I_{n}|=Lx_{n}^{-s}\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L}C_{1}^{-s}\langle c_{n}\rangle^{-s}.$
@@
(2.16) $|\Omega\cap I_{n}|\geq 2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{-1}L^{\prime}}C_{1}^{s}{\gamma^{\prime}}^{\langle c_{n}\rangle^{\tau}}|I_{n}|.$
The inequality (2.7) is obtained from the above by choosing $\gamma$ small enough so that $\gamma<\inf_{n\in\mathbb{N}}(2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{-1}L^{\prime}}C_{1}^{s})^{1/\langle c_{n}\rangle^{\tau}}\gamma^{\prime}$, where the right hand side is positive since $\langle c_{n}\rangle\geq 1$ for all $n\in\mathbb{N}$.
@@
(2.21) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{\prime}>\max\left\{2L\langle L\rangle^{s},\left(\frac{\langle x_{1}\rangle}{x_{0}}\right)^{s}\right\}\geq\max_{k\in\mathbb{N}_{+}}\left\{2L\langle L\rangle^{s},\left(\frac{\langle x_{k+1}\rangle}{x_{k}}\right)^{s}\right\},}$